help!! a rectangle has a length 4 cm greater than it's width.it's area is 45 cm^2.what are the length and width of the rectangle. thanks for helping.
Answer by ankor@dixie-net.com(22740)

a rectangle has a length 4 cm greater than it's width.it's area is 45 cm^2.what are the length and width of the rectangle.
:
Let x = the width of the rectangle
:
It says the length is 4 cm more than the width, therefore we can say:
(x+4) = the length of the rectangle
:
The area is length * width, so we can write it:
x(x+4) = 45
:
x^2 + 4x = 45
:
x^2 + 4x - 45 = 0; a quadratic equation
Factor this to
(x+9)(x-5) = 0
We want the positive solution, when dealing with real objects
x = +5 cm is the width
then
5 + 4 = 9 cm is the length
:
:
Check by finding the area: 9 * 5 = 45
:
:
Did you understand this? Hopefully you can do this type of problems now?
If not, do you have any questions about it?